# Operator Learning with Domain Decomposition for Geometry Generalization in PDE Solving

### v1

@@ A.1 Background on Domain Decomposition
A.2 Revisit on Operator Learning
The goal of operator learning is to learn a mapping $\mathcal{G}:\mathcal{A}\to\mathcal{U}$ between two infinitely dimensional spaces Kovachki et al., (2023). When applied to PDEs, $\mathcal{U}$ is the solution space of a PDE and $\mathcal{A}$ is the space of functions that determine a unique solution of a PDE. Examples of $\mathcal{A}$ are coefficient functions or boundary/initial conditions that defines the PDE and parameters that determine the geometry of domain.
In our study, we decompose any domain into subdomains each of which lives in a distinguished class of basic shapes $\mathcal{P}$. We assume all shapes in $\mathcal{P}$ have Lipschitz boundary and are uniformly bounded, i.e., they are all bounded by a ball $D\subseteq\mathbb{R}^{n}$. We are interested in solving boundary value problems in Eq. (1) in any domain $\Omega\in\mathcal{P}$ with any appropriate boundary condition. We thus separate geometry and boundary conditions from other inputs and represent the input function space of the operator as $\mathcal{A}=\mathcal{P}\times H^{k}(D)\times\mathcal{H}$ where $H^{k}(D)$ is the Sobolev space $W^{k,2}(D)$. The space $\mathcal{P}\times H^{k}(D)$ represents the geometry of the domain together with boundary/initial conditions, $\mathcal{H}$ represents any other input functions such as coefficient function field or source term in the PDE. The neural operator thus approximates the following mapping. Note that in the case of time dependent problem, the space $H^{k}(D;\mathbb{R})$ represents the space of initial condition together with time varying boundary condition and the solution space $\mathcal{U}$ represents a time series up to some time span. The solution operator $\mathcal{G}$ thus has the following form.
@@ A.3 Proof of Theorem 1
Let $u$ and $\tilde{u}$ denote the solution obtained by traditional additive Schwarz method given by Eq. (5) and SNI in Algorithm 1, respectively, with the same initial condition $u^{0}=\tilde{u}^{0}$. Assume $\parallel\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}(\cdot)-\mathcal{S}_{k}(\cdot)\parallel<c$ where $\mathcal{S}_{k}$ is the local solution operator for Eq. 3 for all $k$, and if the traditional algorithm converges, then we have:
  • Convergence: SNI converges to a fixed point;
  • Error bound: there exists a constant $c^{\prime}$ (depending on $c$) such that $\parallel\tilde{u}^{n}-u^{n}\parallel<c^{\prime}$.
@@
(1) Recall that the iterative rules of traditional and neural Schwarz inference are given by:
$u^{n+1}$ $=u^{n}+\tau\sum_{k=1}^{K}\left[R_{k}^{\intercal}\mathcal{S}_{k}(u^{n})-R_{k}^{\intercal}R_{k}u^{n}\right]$ (7)
$\tilde{u}^{n+1}$ $=\tilde{u}^{n}+\tau\sum_{k=1}^{K}\left[R_{k}^{\intercal}\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-R_{k}^{\intercal}R_{k}\tilde{u}^{n}\right]$
where $\tilde{\mathcal{G}}_{k}^{\dagger}$ denotes $\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}$. By simple calculation, we first express the operator $\mathcal{S}_{k}$ explicitly by introducing finite element space under weak formulation: $\mathcal{S}_{k}:u^{n}\mapsto A_{k}^{-1}R_{k}(f-Au^{n})+R_{k}u^{n}$. Hence the convergence of traditional Schwarz algorithm implies $\rho(I-\tau MA)<1$ where $M=\sum_{k=1}^{K}R_{k}^{\intercal}\tilde{A}_{k}^{-1}R_{k}$, and $\rho(\cdot)$ denotes the spectral radius. Hence we have
$\tilde{u}^{n+1}=(I-\tau MA)\tilde{u}^{n}+\tau Mf+\tau\sum_{k=1}^{K}R_{k}^{\intercal}(\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n}))$
Since $\tau Mf$ is a constant term and $\parallel\tilde{\mathcal{G}}_{k}^{\dagger}(\cdot)-\mathcal{S}_{k}(\cdot)\parallel<c$, by random fixed point theorem, the neural algorithm converges to a random fixed point.
(2) Subtraction of Eq. (7) gives:
$\left\|\tilde{u}^{n+1}-{u}^{n+1}\right\|$ $=\left\|(I-\tau\sum_{k=1}^{K}R_{k}^{\intercal}R^{k})(\tilde{u}^{n}-{u}^{n})-\tau\sum_{k=1}^{K}R_{k}^{\intercal}(\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(u^{n}))\right\|$
$=\left\|(I-\tau\sum_{k=1}^{K}R_{k}^{\intercal}R^{k})(\tilde{u}^{n}-{u}^{n})-\tau\sum_{k=1}^{K}R_{k}^{\intercal}(\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n})+\mathcal{S}_{k}(\tilde{u}^{n})-\mathcal{S}_{k}(u^{n}))\right\|$
$\leq(I-\tau MA)\left\|\tilde{u}^{n}-{u}^{n}\right\|+\tau\sum_{k=1}^{K}R_{k}^{\intercal}\left\|\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n})\right\|$
It is easy to see that $\tau\sum_{k=1}^{K}R_{k}^{\intercal}\left\|\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n})\right\|\leq\tau tc$ where $t$ denotes the maximal number of overlapping subdomains. Let $\rho=\rho(I-\tau MA)<1$, we have $\left\|\tilde{u}^{n}-{u}^{n}\right\|\leq\frac{1-\rho^{n}}{1-\rho}\tau tc$. Taking $c^{\prime}=\frac{\tau tc}{1-\rho}$ completes the proof. ∎
@@
Consider the exact operator $A_{k}^{-1}$ and inexact neural operator $\tilde{A}_{k}^{-1}$, $k=1,\cdots,K$. Let $u^{n}$ and $\tilde{u}^{n}$ represent the solutions updated by $A_{k}^{-1}$ and $\tilde{A}_{k}^{-1}$ respectively at the $n$-th step, where the updating rule is given by Eq. (5) with $\tau=1$ and both sharing the same initialization. Suppose that $\parallel A_{k}^{-1}-\tilde{A}_{k}^{-1}\parallel<c$, for $k=1,\cdots,K$, and $\rho(I-MA)<1$, where $M=\sum_{k=1}^{K}R_{k}^{\intercal}\tilde{A}_{k}^{-1}R_{k}$, then we have:
@@
Note that the condition $\rho(I-MA)<1$ is generally challenging to satisfy. To address this issue, we employ the Richardson iteration trick Richardson, (1911) in order to ensure the convergence of the proposed algorithm (Algorithm 1).
A.4 Time-Dependent Problems.
@@
where $\mathcal{L}$ is again self-adjoint and coercive elliptic operator. The additive Schwarz method can be naturally extended to a space-time additive Schwarz method Li and Cai, (2015) by considering a decomposition of the space-time domain $\Omega\times[0,T]$ by taking the product of overlapping decomposition of $\Omega$ and $[0,T]$ respectively. The space-time domain decomposition has the form $\Omega_{i}\times[t_{j-1}-\delta_{T},t_{j}+\delta_{T}]$ where $\delta_{T}$ is the temporal depth and represent overlap in time domain. Once such a decomposition is constructed, the same additive Schwarz method can be applied to the space-time decomposition to get a global solution on the space-time domain, allowing parallel iteration in both space and time domain. Local problems for the above decomposition are again of the form in Eq. (8).
@@
A.5 Symmetries of PDEs
The symmetry group of a general partial differential operator $\mathcal{L}$ refers to a set of transformations that map a solution to another solution, forming a mathematical group. Lie point symmetry is a subgroup of the symmetry group that has a Lie group structure and acts on functions pointwise as transformations on coordinates and function values Brandstetter et al., (2022). In this work, we will in addition be concerned with not just a single operator $\mathcal{L}$, but a family of operators depending on various coefficient fields (e.g., Darcy flow) and various boundary/initial conditions. Symmetries have to be properly extended to these input functions so that a solution with an input function is transformed to another solution with a different input function.
Leveraging these symmetries allows for the generation of an infinite number of new solutions based on a given solution. The idea of utilizing these symmetries as a data augmentation technique for operator learning was initially introduced in Brandstetter et al., (2022). However, we apply these data augmentation to solutions on basic shapes in training local operator and this usage of symmetries echos a point mentioned in (Brandstetter et al.,, 2022, Section 3.2) where the authors point out that these data augmentation can be applied on local patches of solutions instead of the solution on the entire domain.
@@ A.5 Symmetries of PDEs
Equation Lap2d-D Lap2d-M Darcy2d Heat2d
@@
A.6 Discussions
@@
Other formulations of DDMs. Schwarz hybrid formulation discussed in this work is one of the most elementary formulation in DDMs. There are many other more advanced DDMs Mathew, (2008). Steklov-Poincaré framework is based on non-overlapping decomposition and transmission condition as coupling condition for local problems. Langrange multiplier framework leads to the well-known FETI method and is also based on non-overlapping decomposition.
@@
Future Works. Based on the above discussion, there are many potential directions for future works. First, it would be interesting to implement this framework using a message-passing framework instead of an iterative algorithm to accelerate the convergence. Second, extending our framework to address higher-dimensional problems is important, particularly since industrial problems often involve 3-d simulation. Third, more advanced DDMs such as Neumann-Neumann, BDDC and FETI Mathew, (2008) may also be explored. Lastly, other types of PDEs such as saddle point problems and non-linear equations such as Navier-Stokes equation is out of the scope of our current work, and present unique challenge. Tackling these challenges requires not only expertise on operator learning, but also deep understanding of PDEs themselves. We speculate that it would be fruitful to combine rich literatures of DDMs with operator learning.
A.7 Experiments
A.7.1 Datasets
@@
Here we introduce more details of our datasets in both training and testing stage. For training data, we generate random simple polygons with $3\leq n\leq 12$ vertices within $[-0.5,0.5]^{2}$ and create uniform mesh using Gmsh Geuzaine and Remacle, (2008). We prepare a separate dataset for validation during training. For testing data, we generate the three domains depicted in Figure 2 together with mesh using the Gmsh UI. We argue that the complexity of geometric domains is fundamentally determined by their underlying topological and geometrical properties. Based on this intuition, we considered three domains of increasing complexity for evaluation: (1) Domain A: This domain is simply connected, representing the simplest class of geometries; (2) Domain B: This domain has two holes and is multiply connected, indicating a higher level of complexity compared to the simply connected Domain A; (3) Domain C: This domain has one hole with corners, further increasing the geometrical complexity compared to the previous two domains. Through a systematical evaluation across this spectrum of domains, from the simple geometry to more intricate multiply connected domains with holes and corners, we believe the results provide a comprehensive understanding of our framework’s capabilities.
Once the geometries and meshes are created, we specify boundary/initial conditions for various equations and domains and generate solutions using FEniCSx Baratta et al., (2023); Scroggs et al., 2022a ; Scroggs et al., 2022b , a popular open-source platform for solving PDEs with the finite element method (FEM). We adopt Lagrange element of order $1$ (linear element) as our finite element space in generating boundary/initial conditions and solutions. Next we give details on how these boundary/initial conditions and solutions are generated for each type of PDEs. We also summarize these details in Table 5 and 6.
@@
For training data, we discretize the time domain with a fixed time step $t_{s}=0.01$, generate piecewise linear initials and time-varying boundary conditions with values randomly generated within $[0,1]$. $\alpha$ is a random number within $[0.8,1]$. We adopt the backward Euler method Langtangen and Logg, (2017) and generate a time series of $10$ time steps. During training we separate these $10$ times steps into $2$ times series of $5$ times steps and training the neural operator to predict $5$ time steps.
@@
Lap2d-D Lap2d-M Darcy2d Heat2d
@@
Time steps $5$
Inference (SNI) A Partition $K$ 20 20×$\times 16$

### v2

@@ Appendix A Background on Domain Decomposition
Appendix B Revisit on Operator Learning
The goal of operator learning is to learn a mapping $\mathcal{G}:\mathcal{A}\to\mathcal{U}$ between two infinitely dimensional spaces (Kovachki et al., 2023). When applied to PDEs, $\mathcal{U}$ is the solution space of a PDE and $\mathcal{A}$ is the space of functions that determine a unique solution of a PDE. Examples of $\mathcal{A}$ are coefficient functions or boundary/initial conditions that defines the PDE and parameters that determine the geometry of domain.
In our study, we decompose any domain into subdomains each of which lives in a distinguished class of basic shapes $\mathcal{P}$. We assume all shapes in $\mathcal{P}$ have Lipschitz boundary and are uniformly bounded, i.e., they are all bounded by a ball $D\subseteq\mathbb{R}^{n}$. We are interested in solving boundary value problems in Eq. 1 in any domain $\Omega\in\mathcal{P}$ with any appropriate boundary condition. We thus separate geometry and boundary conditions from other inputs and represent the input function space of the operator as $\mathcal{A}=\mathcal{P}\times H^{k}(D)\times\mathcal{H}$ where $H^{k}(D)$ is the Sobolev space $W^{k,2}(D)$. The space $\mathcal{P}\times H^{k}(D)$ represents the geometry of the domain together with boundary/initial conditions, $\mathcal{H}$ represents any other input functions such as coefficient function field or source term in the PDE. The neural operator thus approximates the following mapping. Note that in the case of time dependent problem, the space $H^{k}(D)$ represents the space of initial condition together with time varying boundary condition and the solution space $\mathcal{U}$ represents a time series up to some time span. The solution operator $\mathcal{G}$ thus has the following form.
@@ Appendix C Proof of Theorem 1
Assume the operator $\mathcal{L}$ in Eq. 1 is self-adjoint and coercive elliptic (Mathew, 2008). Let $u^{n}$ and $\tilde{u}^{n}$ denote the iterates of the classical additive Schwarz-Richardson iteration (Eq. 5) and SNI (Algorithm 1), respectively, with the same initialization $u^{0}=\tilde{u}^{0}$.
Let $\mathcal{S}_{k}$ denote the exact local solution operator on $\Omega_{k}$, and define the learned (pre/post-processed) local solver $\tilde{\mathcal{S}}_{k}\triangleq\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}$. Assume the classical iteration mapping is contractive in a norm $\|\cdot\|$ with rate $0<\rho<1$, and the uniform local error bound $\|\tilde{\mathcal{S}}_{k}(\cdot)-\mathcal{S}_{k}(\cdot)\|\leq c$ holds for all $k$. Let $t$ be the maximum number of subdomains covering any degree of freedom. Then we have:
  • Convergence: SNI converges to a fixed point $\tilde{u}^{*}$.
  • Error bound: for all $n$, $\|\tilde{u}^{n}-u^{n}\|\leq c^{\prime}$, and $\|\tilde{u}^{*}-u^{*}\|\leq c^{\prime}$, where one may take $c^{\prime}=\frac{\tau t}{1-\rho}c$.
@@
(1) Write the classical and learned iterations in the same additive Schwarz form:
$u^{n+1}$ $=u^{n}+\tau\sum_{k=1}^{K}R_{k}^{\intercal}\big(\mathcal{S}_{k}(u^{n})-R_{k}u^{n}\big),$ (7)
$ilde{u}^{n+1}$ $=\tilde{u}^{n}+\tau\sum_{k=1}^{K}R_{k}^{\intercal}\big(\tilde{\mathcal{S}}_{k}(\tilde{u}^{n})-R_{k}\tilde{u}^{n}\big),$
where $\tilde{\mathcal{S}}_{k}\triangleq\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}$.
(2) Let $e^{n}\triangleq\tilde{u}^{n}-u^{n}$. Subtracting the two updates in Eq. 7 yields
$e^{n+1}=\big(F(\tilde{u}^{n})-F(u^{n})\big)+\tau\sum_{k=1}^{K}R_{k}^{\intercal}\big(\tilde{\mathcal{S}}_{k}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n})\big),$
where $F(u)\triangleq u+\tau\sum_{k=1}^{K}R_{k}^{\intercal}(\mathcal{S}_{k}(u)-R_{k}u)$ is the classical iteration mapping. By the contractivity assumption of the classical iteration, $\|F(\tilde{u}^{n})-F(u^{n})\|\leq\rho\|e^{n}\|$. Moreover, using the uniform local approximation error bound and the overlap factor $t$ (the maximum number of subdomains covering any degree of freedom), we have
$\left\|\tau\sum_{k=1}^{K}R_{k}^{\intercal}\big(\tilde{\mathcal{S}}_{k}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n})\big)\right\|\leq\tau tc.$
Therefore,
$\|e^{n+1}\|\leq\rho\|e^{n}\|+\tau tc.$
Iterating the inequality gives $\|e^{n}\|\leq\frac{1-\rho^{n}}{1-\rho}\,\tau tc$. Taking $c^{\prime}\triangleq\frac{\tau tc}{1-\rho}$ yields $\|\tilde{u}^{n}-u^{n}\|\leq c^{\prime}$ for all $n$.
(3) Since $u^{n}$ converges by assumption, it is Cauchy. For any $m>n$, we have
$\|\tilde{u}^{m}-\tilde{u}^{n}\|\leq\|u^{m}-u^{n}\|+\|\tilde{u}^{m}-u^{m}\|+\|\tilde{u}^{n}-u^{n}\|\leq\|u^{m}-u^{n}\|+2c^{\prime}.$
Hence $\tilde{u}^{n}$ is also Cauchy and thus converges to some $\tilde{u}^{*}$. Passing to the limit in $\|\tilde{u}^{n}-u^{n}\|\leq c^{\prime}$ yields $\|\tilde{u}^{*}-u^{*}\|\leq c^{\prime}$, completing the proof. ∎
@@
Consider the exact operator $A_{k}^{-1}$ and inexact neural operator $\tilde{A}_{k}^{-1}$, $k=1,\cdots,K$. Let $u^{n}$ and $\tilde{u}^{n}$ represent the solutions updated by $A_{k}^{-1}$ and $\tilde{A}_{k}^{-1}$ respectively at the $n$-th step, where the updating rule is given by Eq. 5 with $\tau=1$ and both sharing the same initialization. Suppose that $\parallel A_{k}^{-1}-\tilde{A}_{k}^{-1}\parallel<c$, for $k=1,\cdots,K$, and $\rho(I-MA)<1$, where $M=\sum_{k=1}^{K}R_{k}^{\intercal}\tilde{A}_{k}^{-1}R_{k}$, then we have:
@@
Note that the condition $\rho(I-MA)<1$ is generally challenging to satisfy. To address this issue, we employ the Richardson iteration trick (Richardson, 1911) in order to ensure the convergence of the proposed algorithm (Algorithm 1).
Appendix D Time-Dependent Problems.
@@
where $\mathcal{L}$ is again self-adjoint and coercive elliptic operator. The additive Schwarz method can be naturally extended to a space-time additive Schwarz method (Li and Cai, 2015) by considering a decomposition of the space-time domain $\Omega\times[0,T]$ by taking the product of overlapping decomposition of $\Omega$ and $[0,T]$ respectively. The space-time domain decomposition has the form $\Omega_{i}\times[t_{j-1}-\delta_{T},t_{j}+\delta_{T}]$ where $\delta_{T}$ is the temporal depth and represent overlap in time domain. Once such a decomposition is constructed, the same additive Schwarz method can be applied to the space-time decomposition to get a global solution on the space-time domain, allowing parallel iteration in both space and time domain. Local problems for the above decomposition are again of the form in Eq. 8.
@@
Appendix E Symmetries of PDEs
The symmetry group of a general partial differential operator $\mathcal{L}$ refers to a set of transformations that map a solution to another solution, forming a mathematical group. Lie point symmetry is a subgroup of the symmetry group that has a Lie group structure and acts on functions pointwise as transformations on coordinates and function values (Brandstetter et al., 2022). In this work, we will in addition be concerned with not just a single operator $\mathcal{L}$, but a family of operators depending on various coefficient fields (e.g., Darcy flow) and various boundary/initial conditions. Symmetries have to be properly extended to these input functions so that a solution with an input function is transformed to another solution with a different input function.
Leveraging these symmetries allows for the generation of an infinite number of new solutions based on a given solution. The idea of utilizing these symmetries as a data augmentation technique for operator learning was initially introduced in Brandstetter et al. (2022). However, we apply these data augmentation to solutions on basic shapes in training local operator and this usage of symmetries echos a point mentioned in Brandstetter et al. (2022, Section 3.2) where the authors point out that these data augmentation can be applied on local patches of solutions instead of the solution on the entire domain.
@@ Appendix E Symmetries of PDEs
Equation Lap2d-D Lap2d-M Darcy2d Heat2d NonLap2d
@@
Appendix F Time Complexity
F.1 Empirical Time Complexity
We provide empirical results on runtime for our main results. We discuss how better initialization can accelerate the whole iterative process in the next point.
We first provide empirical runtime on a single sample for each stationary problem and each domain reported in our main result. We use the following metrics:
  • Time to convergence (TTC).
  • Time to 15/10/5% relative $l_{2}$ error following the practice in Mao et al. (2024).
Equations Domains TTC(s) TT15%(s) TT10%(s) TT5%(s)
Laplace2d-Dirichlet A 100 40 50 70
Laplace2d-Dirichlet B 269 107 137 194
Laplace2d-Dirichlet C 28 8 11 17
Laplace2d-Mixed A 162 82 103 137
Laplace2d-Mixed B 714 511 620 -
Laplace2d-Mixed C 68 43 52 -
Darcy2d A 84 37 54 -
Darcy2d B 247 144 176 -
Darcy2d C 26 12 15 -
Table 4: Empirical runtime for different equations and domains
Factors that affect the runtime are:
  1. 1.
    Type of equations. We observe that Laplace2d-Mixed takes longer on all domains. We also observe that the existence of Neumann boundary condition leads to a larger range of function values for the solution of Laplace2d-Mixed. This leads to more iterations steps required to reach convergence.
  2. 2.
    Number of subdomains $K$ and step size $\tau$. In the above table, domain B takes longer for all equations because it has 40 subdomains compared to 20 for A and C. A large number of subdomains leads to more time consumption for an iteration. We illustrated how choice of $\tau$ affects the number of iterations to convergence in section 4.3 of our paper.
  3. 3.
    Local operator architecture. While GNOT gets better results in accuracy, a drawback of transformer-based methods is that they are usually slower than FNO (Hao et al., 2023).
  4. 4.
    Initialization. This is discussed in the next point.
We note that our implementation is not optimized to fully parallelize the iterative process; for example, the normalization process is not parallelized in our implementation.
Numerical solvers are very fast in generating solutions for the domains we tested on and we do not expect our approach to be faster than these highly optimized numerical solvers on these (still) simple domains. As a reference, generating a solution for Laplace2d-Dirichlet on domain A using classical FEM solution takes $6.15\times 10^{-4}$ seconds and performing a GNOT inference on the same domain takes $1.26\times 10^{-2}$ seconds. We can see that even classical numerical solver is faster than GNOT inference. However, DDMs are a conventional approach implemented in commercial software designed to solve PDEs on large-scale and complicated domains. We replace the local FEM solver in DDMs by a data-driven neural operator and thus expect our approach to show superiority when the problem domain is large and complicated.
F.2 Acceleration through Better Initialization
We discuss how to accelerate the iterative process by starting with a better initialization. In the original implementation, we always start with a zero solution in the interior of the domain. To accelerate the process, we initialize with solutions from GNOT direct inference and find that it considerably saves our time. We report the time consumption on Laplace2d-Dirichlet using the same metrics as the previous table. The only difference is in initialization.
Equations Domains TTC(s) TT15%(s) TT10%(s) TT5%(s)
Laplace2d-Dirichlet A 28 8 11 17
Laplace2d-Dirichlet B 70 3 6 21
Laplace2d-Dirichlet C 20 1 2 7
Table 5: Runtime results with improved initialization for Laplace2d-Dirichlet equations
However, coming up with a better initialization is not trivial and can be an interesting future work.
Appendix G Discussions
@@
The fundamental "local-to-global" principle of our method remains unchanged: decompose a complex 3D domain into smaller, canonical 3D subdomains, solve locally with a neural operator, and iteratively stitch the solutions. However, implementation faces additional challenges in 3D. For example, generating building blocks in 3D. One way is to use polyhedrons as building blocks in 3D. Another choice is to generate building blocks from "growing" tetrahedra in 3D. Both can be very interesting ways to explore. In summary, extending our method to 3D is a challenging but highly promising research trajectory. It is not a simple "plug-and-play" extension but requires careful research at the intersection of computational geometry, operator learning, and high-performance computing.
Other formulations of DDMs. Schwarz hybrid formulation discussed in this work is one of the most elementary formulation in DDMs. There are many other more advanced DDMs (Mathew, 2008). Steklov-Poincaré framework is based on non-overlapping decomposition and transmission condition as coupling condition for local problems. Langrange multiplier framework leads to the well-known FETI method and is also based on non-overlapping decomposition.
@@
Future Works. Based on the above discussion, there are many potential directions for future works. First, it would be interesting to implement this framework using a message-passing framework instead of an iterative algorithm to accelerate the convergence. Second, extending our framework to address higher-dimensional problems is important, particularly since industrial problems often involve 3-d simulation. Third, more advanced DDMs such as Neumann-Neumann, BDDC and FETI (Mathew, 2008) may also be explored. Lastly, other types of PDEs such as saddle point problems and non-linear equations such as Navier-Stokes equation is out of the scope of our current work, and present unique challenge. Tackling these challenges requires not only expertise on operator learning, but also deep understanding of PDEs themselves. We speculate that it would be fruitful to combine rich literatures of DDMs with operator learning.
Appendix H Experiments
H.1 Datasets
@@
Here we introduce more details of our datasets in both training and testing stage. For training data, we generate random simple polygons with $3\leq n\leq 12$ vertices within $0.5,0.5]^{2}$ and create uniform mesh using Gmsh (Geuzaine and Remacle, 2008). We prepare a separate dataset for validation during training. For testing data, we generate the three domains depicted in Figure 2 together with mesh using the Gmsh UI. We argue that the complexity of geometric domains is fundamentally determined by their underlying topological and geometrical properties. Based on this intuition, we considered three domains of increasing complexity for evaluation: (1) Domain A: This domain is simply connected, representing the simplest class of geometries; (2) Domain B: This domain has two holes and is multiply connected, indicating a higher level of complexity compared to the simply connected Domain A; (3) Domain C: This domain has one hole with corners, further increasing the geometrical complexity compared to the previous two domains. Through a systematical evaluation across this spectrum of domains, from the simple geometry to more intricate multiply connected domains with holes and corners, we believe the results provide a comprehensive understanding of our framework’s capabilities.
Once the geometries and meshes are created, we specify boundary/initial conditions for various equations and domains and generate solutions using FEniCSx (Baratta et al., 2023; Scroggs et al., 2022a; b), a popular open-source platform for solving PDEs with the finite element method (FEM). We adopt Lagrange element of order $1$ (linear element) as our finite element space in generating boundary/initial conditions and solutions. Next we give details on how these boundary/initial conditions and solutions are generated for each type of PDEs. We also summarize these details in Table 7 and 8.
@@
For training data, we discretize the time domain with a fixed time step $t_{s}=0.01$, generate piecewise linear initials and time-varying boundary conditions with values randomly generated within $[0,1]$. $\alpha$ is a random number within $[0.8,1]$. We adopt the backward Euler method (Langtangen and Logg, 2017) and generate a time series of $10$ time steps. During training we separate these $10$ time steps into $2$ time series of $5$ times steps and training the neural operator to predict $5$ time steps.
@@
NonlinearLaplace2d. A nonlinear Laplace equation in 2d with pure Dirichlet boundary condition following an example in Langtangen and Logg (2017). The governing equation is
$\nabla\cdot((u^{2}+1)\nabla u)$ $=0\quad\quad\text{in}\ \ \Omega$ (13)
$u$ $=u_{D}\quad\ \text{on}\ \ \partial\Omega.$
Lap2d-D Lap2d-M Darcy2d Heat2d NonlinearLap2d
@@
Rot.
Time steps 55
Inference (SNI) A Partition $K$ 20 $\times 16 20$